You are given the following information about a closed economy with no government: Consumption = 115 + 0.6Y Investment = 550 Use the above information to answer the questions that follow:
Q.4.2 Calculate the value of the multiplier
1
Expert's answer
2020-09-22T14:01:35-0400

Solution:

The value of the multiplier = "\frac{1}{(1-MPC) }"


MPC = "\frac{Change \;in \;Consumption\;(C) }{Change\;in \;Income(Y) }"


"Y = C+I"

"Y = 115+0.6Y+550"

"Y = 115+550+0.6Y"

"Y = 665+0.6Y"

"Y-0.6Y=665"

"0.4Y=665"

"Y=\frac{665}{0.4}"


"Y = 1663"


"C=115+0.6Y"

"C=115+0.6(1663)"

"C=115+998"

"C=1113"


"MPC=\frac{1113}{1663}" = "0.7"


"Multiplier =\frac{1}{(1-MPC) }"


"Multiplier =\frac{1}{(1-0.7) }"


"Multiplier = \frac{1}{0.3}"


"Value\;of \;the\; multiplier = 3.33"